OpenStudy (anonymous):

How long, in years, will it take 80 kg of plutonium (half life 24 000 years) to decay to approximately 1.8 kg?

OpenStudy (anonymous):

solve \[1.8=80\times (\frac{1}{2})^{\frac{t}{24,000}}\] for \(t\)

OpenStudy (anonymous):

first divide both sides by 80 to get \[.0225=(\frac{1}{2})^{\frac{t}{24,000}}\] then change of base gives \[\frac{t}{24,000}=\frac{\ln(.0225)}{\ln(.5)}\] and multiply by 24,000 and use a calculator

OpenStudy (anonymous):

okay and can u see if i have the right answer for this?

OpenStudy (anonymous):

i got 228 years

OpenStudy (anonymous):

no that cannot be right

OpenStudy (anonymous):

if the half life is 24,000 years then that is the time it will take to go from 80 to 40

OpenStudy (anonymous):

oh sorry this answer is for a different questionj and i wanted u to check if i got the answer right

OpenStudy (anonymous):

looks like about 131317 http://www.wolframalpha.com/input/?i=24000*ln%28.0225%29%2Fln%28.5%29

OpenStudy (anonymous):

260 grams of a radioactive substance decays to 160 grams after 160 years. To the nearest year, what is the half-life of the substance?

OpenStudy (anonymous):

is the answer for this 228

OpenStudy (anonymous):

\[160=260(\frac{1}{2})^t\] \[\frac{8}{13}=(\frac{1}{2})^t\] \[t=\frac{\ln(\frac{8}{13})}{\ln(.5)}\]

OpenStudy (anonymous):

sorry if im asking too many questions. can u help me out with a few more if you dont mind?

OpenStudy (anonymous):

damn i made a mistake. have to multiply my answer by 160

OpenStudy (anonymous):

i get 112

OpenStudy (anonymous):

so what ever you get for t you multiply it by 160?

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

equation at the beginning should have been \[160=260\times (\frac{1}{2})^{\frac{t}{160}}\]

OpenStudy (anonymous):

can u help me with this one? A stock that Jane bought at $75 a share is now worth $629.86 a share. If it was averaged a 12% increase in value a year, then how many years has Jane owned the stock? Round off your answer to one decimal place.

OpenStudy (anonymous):

solution method is identical, but you should end up with \[\frac{t}{160}=\frac{\ln(\frac{8}{13})}{\ln(.5)}\] \[t=160\times \frac{\ln(\frac{8}{13})}{\ln(.5)}\]\]

OpenStudy (anonymous):

\[629.86=75(1.12)^t\] solve for t

OpenStudy (anonymous):

so would it be ln629.86/ ln 75

OpenStudy (anonymous):

?